### 17 Mar 2026 v3

# Interferometric discrepancy between the non-relativistic solution to the Klein-Gordon and Schrödinger wave equations due to their dissimilar phase velocities

Abstract
Adding a constant energy offset leaves classical dynamics unchanged. In quantum mechanics it changes the phase velocity of the wavefunction. The inclusion of the constant rest energy in the Klein-Gordon formulation leads to significantly higher phase velocities compared with the Schrödinger equation. The Schrödinger equation predicts an attenuation of the wavefunction along one of the paths in a Sagnac interferometer when a beamsplitter’s trajectory along that path includes a segment where its speed exceeds the phase velocity of a free particle. Such an attenuation does not occur for electromagnetic waves nor for eigenstates of momentum in the Klein-Gordon equation since the speed of the beamsplitter cannot then exceed the phase velocity of the wave. This attenuation reduces the amplitude without introducing a phase shift, preserving the overall structure of the transmitted wave group. While a Klein-Gordon wave group undergoes three traversals of the beamsplitter that moved, it experiences attenuation equivalent to only a single pass, whereas the Schrödinger equation predicts the expected attenuation for three passes. This discrepancy highlights a fundamental incompatibility between the Klein-Gordon equation and the Schrödinger equation in a regime where their predictions should converge.
pacs:
03.65.Pm, 07.60.Ly
Significance statement
In classical mechanics, adding a constant offset to the energy does not change the equations of motion. In quantum mechanics, the phase velocity (speed of wave crests in the particle’s wavefunction) varies with such an addition. The two theories that describe non-relativistic quantum mechanics should predict the same results but have vastly different phase velocities; one adds a constant rest energy, $mc^{2}$, while the other, the Schrödinger equation, does not. The latter allows a beamsplitter to outpace the wave crests, inducing wave function attenuation not possible in the former. These differing predictions restrict the validity of either adding a rest energy term (non-relativistic Klein-Gordon equation) or not (Schrödinger equation).
I Introduction
The phase velocity of a free particle plane wave solution (momentum eigenstate) of the Schrödinger equation is $\omega/k=KE/p=v_{g}/2$ where $KE=p^{2}/2m,p$, and $v_{g}$ are the kinetic energy, momentum, and group or particle velocity. In the non-relativistic limit the phase velocity for eigenstates of momentum in the Klein-Gordon equation is, in this case, $\omega/k=(mc^{2}+KE)/p\approx c^{2}/v_{g}$ where $c$ is the speed of light.
This difference in phase velocities arises from the rest energy term present in the Klein-Gordon eqn. but not in the Schrödinger eqn. Using the ansatz $\Phi=\exp[-imc^{2}t/\hbar]\Psi$ in the Klein-Gordon eqn. and the assumption that the free particle kinetic energy is much less than the rest energy results in $\Psi$ satisfying the Schrödinger eqn. drell The frequency of the free particle wavefunction in the Klein-Gordon eqn. is then $(mc^{2}+mv_{g}^{2}/2)/\hbar$.
Solutions to the Dirac eqn. also satisfy the Klein-Gordon eqn. for each of its four components. Interferometry is governed by these eqns. and has been observed, in the non-relativistic regime, for electrons, neutrons, atoms, and molecules.cronin
All of the Fourier components of a Gaussian wave group then contain the global phase factor, $\exp[-imc^{2}t/\hbar$]. Since the probability density function (PDF) is defined as $|\Psi|^{2}$, this factor vanishes, yielding the Schrödinger result. Although both have the same structure, wave crests move through the wave group at vastly different speeds in these two cases.
At the output port of the interferometer described below the Schrödinger and Klein-Gordon PDFs are equivalent in structure. However, their differing phase velocities lead to dissimilar attenuation when interacting with a moving beam splitter.
It is often remarked that the phase velocity of an eigenstate of momentum in quantum mechanics is not a physically meaningful quantity since it can exceed the speed of light and carries no information. However, the phase difference between two paths is measurable, generating interference that is evident in the PDF at the output port of an interferometer. If the size of the wave group traversing different paths is larger than this path difference then the interference is determined by the central momentum of the wave group, essentially matching the response of an eigenstate with that momentum. The following analysis begins with this case and later extends to wave groups.
The significance of the phase velocity in interferometry is often concealed by using a snapshot method for determining the difference in phase; time is fixed while the number of wave crests is compared along the two paths. The retarded phase method, however, emphasizes the importance of the phase velocity in such calculations.kowalski2010
For example, let the phase at the input port of an interferometer be given by $\omega_{0}t$. The phase at the output port along one path is then $\omega_{0}t_{ret}$, where $t_{ret}$ is the retarded time for that path (the time at which the wave crest at the output port was at the input port). For light traversing a static interferometer $t_{ret}=t-t_{out}$, where the wave crest transit time to the output port is $t_{out}=L/c$ and $L$ is the distance that it travelled.
One advantage of this method is that the phase along time dependent path lengths is more easily determined. An example is a path that involves reflection of a particle in an eigenstate of momentum from an accelerating mirror. The reflected states for the Schrödinger and the Klein-Gordon eqns. then differ in the non-relativistic limit due to the difference in phase velocities (although they don’t differ for constant mirror velocity).kowalski2010
Both the snapshot and retarded phase methods can also be applied to the electric Aharonov-Bohm effect where one segment of the particle’s trajectory lies within a field-free region characterized by a constant potential. The Schrödinger Hamiltonian then includes a constant term that modifies the phase velocity $\omega/k=(qV+KE)/p=qV/p+v_{g}/2$. The phase velocity for the Klein-Gordon eqn. in this case is $\omega/k=(mc^{2}+qV+KE)/p=c^{2}/v_{g}+qV/mv_{g}+v_{g}/2$. However, the phase velocity is now dependent on the sign of $V$, enabling it to be increased or decreased and, in principle, counteracting the effect of the rest energy term. An optical interferometer employs a dielectric that changes the phase velocity along a segment of the path but does not change the frequency of the wave (however, the frequency does change when a neutron enters moving matter horne ).
In the electric Aharonov-Bohm effect, a difference in electrostatic potential along two paths induces a measurable quantum phase shift—a phenomenon not produced by a constant rest energy offset. However, the following interferometer produces amplitude-attenuated, rather than phase-modulated, interference, with attenuation determined by the phase velocity.
II Results
II.1 Plane-wave interference
The retarded phase method can reveal interferometry that might be non-intuitive when considering only the snapshot method. An example for the Schrödinger equation is the one-dimensional nontrivial motion of a beamsplitter, BS, that exceeds the phase velocity on part of its trajectory, as shown in the lab frame schematically in fig. 1. The linear motion of this BS (not the stationary input-output port beamsplitter for the Sagnac interferometer) is constrained to be within the Sagnac interferometer. The BS is initially treated as an ideal (zero-thickness, achromatic) element, with more realistic models presented later. The BS is initially stationary as shown in fig. 1(a) then moves faster than the speed of a wave crest in fig. 1(b) before coming to rest in fig. 1(c).
The eigenstate of momentum injected into this interferometer is represented only by segments for clarity. For example, the two largest amplitude wave crest segments shown in fig. 1(a), labeled $A$ and $B$, represent the clockwise (CW) and counterclockwise (CCW) wavefunctions before traversing the BS. The segments labeled $A$ and $B$ with reduced amplitudes represent the waves after traversing the BS. The effect of reflections (not shown in the diagram) from the beam splitter are suppressed by anti-reflection coatings or by tilting the beamsplitter to misalign back-reflections away from the detector.
Figure 1: Interference generated by one dimensional motion of a beamsplitter, BS, within a Sagnac interferometer for the Schrödinger equation. Only segments of the wave trains are shown. The highest amplitude segments in (a) illustrate the counter-propagating waves prior to traversing the BS. The length of the arrows indicate the speeds of the wavecrests and BS while the amplitude of a wave segment indicates how often it traversed the BS. The source, S, emits a harmonic wave that interacts with a stationary BS in frame (a). The BS then moves to the left faster than the CW traveling wave in frame (b) generating a transmitted wave that is left behind it. The BS then comes to rest after having passed through the initially transmitted CW sequence of wave crests (shown in frame(a)) in frame (c). This lagging wave crest sequence then traverses the BS for a final time, reducing its amplitude again. The BS travels a distance $L$.
The CW wave segment overtaken by the BS experiences three traversals of the BS, compared to a single traversal for the CCW segment. Since the difference in path lengths between the CW and CCW paths is zero the particle beam intensity at the output of the Sagnac interferometer, after the BS has come to rest, is given by $I_{Sch}=I_{ccw}+I_{cw}+2T^{2}\sqrt{I_{ccw}}\sqrt{I_{cw}}$, where $T$ is the BS transmission coefficient and $I_{ccw}$ and $I_{ccw}$ are the intensities for one traversal of the BS. The interference of these counter propagating waves with different attenuations lasts only until all the wave crests that were left behind the BS have transmitted through it again when it comes to rest. For no motion of the BS $I_{Sch}^{0}=I_{ccw}+I_{cw}+2\sqrt{I_{ccw}}\sqrt{I_{cw}}$.
No such variation in the interference due to motion of the BS can occur for wave crests of electromagnetic radiation or for those obeying the Klein-Gordon equation since the speed of the BS can never exceed that of a wave crest. The CW and CCW waves then traverse the BS only once for these wave equations.
The only other case of a BS moving faster than wave crests involves wave propagation in material media. For example, consider the acoustic wave equation for the apparatus shown in fig. 1. A shock wave appears in front of the BS as it exceeds the speed of a wave crest (due to the compression in the density of the medium). However, a small acoustic receiver following the trajectory of the BS in fig. 1 will measure the same wave train segment twice: first while stationary, and again after overtaking this wave segment and returning to rest.
The wave reflected from the moving BS is Doppler shifted. However, the transmitted wave is an amplitude-attenuated replica of the incident wave. No shock wave occurs for the Klein-Gordon and Schrödinger equations. However, only the latter’s dispersion relation supports phase velocities much slower than the speed of light.
The quantum behavior of the BS, described above, can be derived using two separate approaches: (1) transformation to and from the rest frame of the BS, or (2) the application of the retarded phase method in the lab frame. The former method is straightforward to analyze since it only involves static boundary conditions. However, this transformation to and from the BS frame for the Schrödinger equation requires a unusual phase factor.levi
During most of its trajectory the BS is either at rest or moving at constant velocity. A small duration, however, involves acceleration, the effects of which are approximated as a sequence of momentarily co-moving inertial frame interactions between the wave and the BS. The calculation utilizing these transformations between the lab and BS frame is outlined next.
Consider first the interference at the output port of the Sagnac interferometer in the lab frame for the Schrödinger equation while the BS moves at constant speed $V$. The BS interacts with two waves in fig. 1: (I) when both the BS and incident wave crests move in the same CW direction and (II) when the moving BS and the CCW wave crests move in opposite directions.
The CW wave crests that move in the same direction as the BS in the lab frame for case (I) now move toward the BS in its frame. Continuity of the incident, transmitted, and reflected wave functions and their first derivatives is trivially satisfied in the BS frame. The results are then transformed back to the lab frame where the transmitted wave crests now move behind the BS in the direction of its motion. The reflected wave crests are not considered since they never reach the detector.
The transmitted wave crests move slower than the BS and are left behind it as shown in fig. 1(b). These lagging wave crests have the same frequency, $mv_{g}^{2}/(2\hbar)$, and wavevector, $mv_{g}/\hbar$, as those of the wave before interacting with the BS and travel in the same direction as the BS. After the BS comes to rest these wave crests transmit through the BS again for a total of three times before reaching the detector.
Consider case (II) in the frame of the BS. The incident wave crests (ccw) move faster toward the BS than in case (I). The reflected crests are again disregarded since they do not enter the detector due to the orientation of the BS. The wave crests transmitted through the BS in the lab frame have the same frequency, $mv_{g}^{2}/(2\hbar)$, and wavevector, $mv_{g}/\hbar$, as before interacting with the BS while moving in the direction opposite to that of the BS. Their amplitude is that associated with one transmission through the BS.
Contrast this calculation with that using the Klein-Gordon equation where the wave crests cannot be overtaken by motion of the BS. Boundary conditions are applied in the BS frame after a Lorentz transformation from the lab frame. After interaction, the wave’s parameters are transformed back to the laboratory frame, where the reflected phase velocity is $c^{2}/(2V+v_{g})$. This result (in addition to that for the same interaction in the Schrödinger equation) can also be obtained using the retarded phase method without the need for a transformation to and from the frame of the BS.kowalski2010 Again, the transmitted wave crests are unaffected by the motion of the BS.
Consider the moving BS in the lab frame for case (I) in the Klein-Gordon equation. The transmitted wavecrests move ahead, rather than lagging behind it (as they do in the Schrödinger eqn.) After the BS comes to rest the CW and CCW waves have traversed the BS only once. The particle beam intensity at the output of the Sagnac interferometer, $I_{KG}=I_{ccw}+I_{cw}+2\sqrt{I_{ccw}}\sqrt{I_{cw}}$, the same as with a stationary BS. This result also applies for electromagnetic, EM, waves.
The crucial property of the moving or stationary BS is its ability to attenuate the wave without affecting its phase. The only information about the interaction that is carried by the transmitted wave is its attenuation. In contrast, the motion of a mirror or the BS can be inferred from the phase shift of the reflected wave.
II.2 Plate and grating beamsplitters
Experimental verification depends on the BS behaving in this manner. A plate or slab of material that is used as a BS illustrates complications that can arise. Retro-reflection occurs at the interfaces, while transmission propagates through the slab’s moving medium; both processes are characterized by an index of refraction. The phase shift of neutrons propagating through such a slab moving at constant speed has been investigated.horne
The effect of accelerating the slab to sub-phase velocity speeds has been calculatedkowalski1993 and verified experimentally.frank The frequency shift of the transmitted wave crests is $\Delta\nu=mgL(1-n)/n\hbar$ where $m$, $g$, $L$, and $n$ are the neutron mass, slab acceleration, slab length, and slab index of refraction, respectively.
This system is studied to examine the validity of the Equivalence Principle, specifically by showing that wave propagation within an accelerated assembly (source, slab, and detector) is indistinguishable from that in a uniform gravitational field. A related experiment involves dropping the slab in a uniform gravitational field. The effect of acceleration is then related to a displacement of the slab in its inertial framebalzas as it falls freely; the loss in potential energy due to this displacement in the gravitational setting results in an increase in frequency of the wave traversing the slab. In a similar setting, slab displacement measurements have been made of order the Planck length when interacting with EM radiation in a gravitational field.kowalski1993b ; kowalski2022
Acceleration of the slab in the direction (or in the opposite direction) of wave crest motion increases (or decreases) the frequency of the transmitted wave. The transmitted wave train then acquires frequency chirps during the acceleration and deceleration stages of the slab trajectory. The frequency of the transmitted wave train during the constant speed part of the trajectory does not change.horne These finite segments of frequency chirps then disperse or spread as they move through the wave train. The frequency of the wave crests (or speed of the transmitted particle during slab acceleration) increases or decreases for $n>1$ or $n<1$, respectively.
A grating beamsplitter avoids these accelerating matter complications since the phases of the incident and zeroth order diffracted beams are the same at the grating apertures. The zeroth order beam is attenuated via loss into higher order diffracted beams. In fig. 1(a) the zeroth order beam transmits through the grating generating the CW wavetrain that is later overtaken by the grating BS in fig. 1(b). After the grating comes to rest the wavetrain left behind it then transmits in zeroth order through the grating again (this wavetrain has then traversed the grating three times) and interferes with the zeroth order CCW beam (that has traversed the grating one time). The grating is oriented so that specularly reflected light is directed away from the detector.
During the acceleration phase, the BS momentarily reaches the same velocity as the particle. At this time, in the frame of the BS, the Schrödinger wavefunction has neither spatial nor temporal dependence while the Klein-Gordon wavefunction has a purely time-dependent harmonic oscillation at the Compton frequency. For BS velocities on both sides of this condition the wavelength is large. The effect of resonance driven dispersion is a reduced index of refraction as the wavelength increases. As higher-order diffracted beams are eliminated in the grating BS, from this increase in wavelength, more energy is redirected into the zeroth-order transmission.
This causes a transient spike in transmission through both BS types during acceleration and deceleration. This system is modeled as a superposition of two wavefunctions: the steady-state transmission through the BS and the transient increase in transmission due to acceleration. Because the latter component travels faster than the phase velocity, it can be separated from the effect illustrated in fig. 1 while the former component is that shown in this fig.
II.3 Wave group interference
Consider next motion of a wave group through the Sagnac interferometer. The group velocity, $d\omega/dk$, for the Schrödinger eqn. is the same as that for Klein-Gordon equation since they differ only by a constant rest energy term.
Let a free particle of mass $m$, that is in a Gaussian wave group state of width $\sigma\ll L$, be launched into the interferometer. $L$ is the distance between the initial and final positions of the grating BS during its trajectory. This BS begins its motion when the CW wave group has just finished traversing the BS and finishes its motion after the BS has overtaken the wave group. The transient transmission spike occurring as the BS accelerates through an eigenstate of momentum is eliminated since now the BS reaches a speed faster than the group and phase velocities before transiting the wave group.
At the interferometer input port the relativistic wavefunction is $\Phi(\mathbf{x},t)=\psi(\mathbf{x},t)e^{-imc^{2}t/\hbar}$, where $\psi(\mathbf{x},t)$ is the slowly varying Gaussian Schrödinger wavefunction. The Klein-Gordon PDF, $|\psi|^{2}$, then matches that of the Schrödinger wavefunction at this input port.
Let the attenuation coefficient of the grating BS be achromatic. The CCW wave group is attenuated once as it traverses this BS but it otherwise moves exactly as if it had not interacted with the BS.
The Fourier components of the CW wave group that have phase velocities greater than that of the BS will only experience one attenuation from the BS. However, those with phase velocities less than the velocity of the BS will experience attenuation three times resulting in a distorted wave group at the interferometer output port. As the BS speed increases, this distortion decreases and is assumed to be negligible.
Each of the momentum eigenstates that form the wave group experiences the same attenuation factor resulting in the same attenuation for the wave group. The wave group interference at the output port for the Schrödinger equation then shifts from $I_{Sch}^{0}$ when the BS remains stationary to $I_{Sch}$ when it follows the trajectory described above.
For the Klein-Gordon eqn. all momentum eigenstates experience only one transit through the BS. The wave group interference is given by $I_{KG}$ and is independent of the BS trajectory.
[…]
The dashed arrows show the phase velocity for the central Fourier component of the wave group (the other components contribute to its variation in amplitude as it moves through the wave group envelope) while the solid arrows indicate the velocity of the wave group. The star symbol just above a particular wave crest follows its trajectory.
[…]
Figure 2: Interaction of a wave group along the BS trajectory of fig. 1, as viewed in the rest frame of the BS.
The Schrödinger equation allows the localized wave group to be interpreted as the particle’s probable position, with the observed attenuation corresponding to the wave group traversing the beam splitter (BS) three times. In the Klein-Gordon equation the wave group traverses the BS three times while the attenuation corresponds to only one pass through the BS. Although the EM wave equation also predicts only one attenuation of the wave group this result occurs because the BS can never overtake a wave group that travels at the speed of light.
III Discussion
Whether a harmonic wave or a wave group is injected into the interferometer, the Schrödinger interferometer output is the same; it changes from $I_{Sch}^{0}$ for a stationary BS to $I_{Sch}$ for the above BS trajectory. The Klein-Gordon output, $I_{KG}$, remains the same for both harmonic waves and wave groups regardless of BS motion.
Fig. 1 (b) shows the BS moving through the CW harmonic wave, producing a decrease of the wavefunction in the transmitted amplitude that is left behind the BS. This diagram omits the reflected wave (or diffracted waves in a grating BS), which, combined with the incident and transmitted waves, satisfies the moving BS boundary conditions. The solution to the Schrödinger eqn. then involves eigenstates of momentum for the incident, reflected, and transmitted waves in front of and behind the BS. The change in amplitude between the incident and transmitted waves is intrinsically linked to the BS and does not represent a discontinuity of the wave function that travels at the particle or group velocity.
In the non-relativistic limit, solutions to the Klein-Gordon eqn. include rest mass energy, whereas the Schrödinger eqn. defines energy solely in terms of kinetic and potential components. Rest energy enables the modeling of energy transfer to internal degrees of freedom during inelastic collisions, while this is not possible with the Schrödinger eqn.kowalskiedu However, the BS only attenuates the wavefunction via reflection/transmission or diffraction. This process can be modeled with the Schrödinger equation since the particle’s energy in not altered.
Interference generated at the output port of this Sagnac interferometer provides no information about the BS trajectory in the EM and Klein-Gordon equations since the PDF is static. However, the PDF for the Schrödinger equation varies from $I_{Sch}^{0}$ to $I_{Sch}$. Information on the BS trajectory is absent; the PDF only indicates that the phase velocity is lower than that of the BS. The total energy of the particle is determined by such a PDF measurement of the the phase velocity.
While fig. 1 depicts interference using macroscopic beamsplitters, the method also applies to microscopic systems. For instance, an atom can act as a beamsplitter through its ability to simultaneously scatter and not scatter a non-zero rest mass body. In this case quantum correlated interference between the particle and BS may become important. Similar interferometric correlations have been studied.kowalskicorrelation
Instead of using the interferometer shown in fig. 1, a direct measurement can be made with the same BS trajectory. In this case, the input-output Sagnac beamsplitter is removed. The wave group from the source then travels only in the CW direction. The detector measures the PDF of the wave group that is attenuated three times in the Schrödinger equation. The Klein-Gordon equation predicts a PDF that has been attenuated by only one pass through the BS, yet the wave group passes three times through the BS.
Interference is not path-dependent in this case. Instead, the wave group’s shape is determined by the interference between its Fourier components. These are uniformly attenuated in the Klein-Gordon equation. In the Schrödinger equation certain components escape threefold attenuation when the BS moves slower than their phase velocity, thereby distorting the PDF. This demonstrates the influence of phase velocity on wave packet dynamics, independent of interferometric methods.
Experimental verification of the interference shown in fig. 1 is difficult with particles of non-zero rest mass. However, neutronkawasaki and atomzhou Sagnac interferometers have been constructed. Neutrons can be generated at speeds of $10\rightarrow 1000$ m/s for cold/thermal beams, and down to $5$ m/s for ultracold neutron beams. With laser cooling the speeds of atoms can be reduced to less than $0.3$ m/s. Surpassing these speed values with a BS is not unreasonable.
In classical mechanics, adding a constant energy offset to the Hamiltonian has no effect on the dynamics of a system. In quantum mechanics, this addition alters the phase velocity and introduces a global phase factor to the Fourier components of the Gaussian wave group. For a static BS the transmitted PDF is the same with and without this constant offset. However, two speeds are involved in fig. 1: the phase velocity and the BS speed. Their relative magnitudes determine the level of attenuation.
Experimental confirmation of the Klein-Gordon prediction would demonstrate that the Schrödinger equation has limited validity, failing in regimes where it was previously expected to hold. The discrepancy between the three-pass wave group traversal of the BS and the resulting single-pass attenuation of the wave group warrants a reevaluation of how the PDF is interpreted using the Born rule with the correspondence principle.
Conflict of interest
The author has no relevant conflicts of interest to disclose.